# Circumbinary disks in post common envelope binary systems with compact objects

### v1

@@ 3.2 The role of the parameter $\beta$
The parameter $\beta$ that we introduced in Section 3.1 can, in principle, take any value from zero to somewhat larger than 1. To demonstrate the role of $\beta$ in determining the number of post-CEE systems that form a CBD, we re-analyze the population synthesis data from Grichener (2023) and examine three final populations that go through the CEE of a giant star with a compact object, and end as NS-NS, BH-BH, and NS-BH binaries.
In Figure 1, we present the percentage of NS-NS systems (blue dots), BH-BH systems (green dots), and NS-BH systems (red dots) that form a post-CEE CBD as a function of the parameter $\beta$. As expected, higher values of $\beta$, which represent systems that retain more of their initial angular momentum, correspond to a greater percentage of systems forming a post-CEE CBD.
Figure 1: The percentage of NS-NS (blue dots), BH-BH (green dots), and NS-BH (red dots) systems that can form a CBD, from all systems that end as NS-NS systems, BH-BH systems, and NS-BH systems, respectively. Statistical error bars are smaller than the marker size.
The graph shows that a CBD is formed in about half of the systems for $\beta=0.197$ in BH-BH binaries, $\beta=0.304$ in NS-NS binaries, and $\beta=0.498$ in NS-BH binaries. The varying percentage of post-CEE CBD formation between the different binary populations for a given value of $\beta$ results from the different values of the ratio of the final orbital radius (semi-major axis) to the initial giant’s radius, the spiraling-in factor $\eta_{\rm SIF}\equiv a_{\rm f}/R_{\ast,i}$ (equation 4). As we will see in Section 4, on average, BH-BH binaries have the smallest values of $\eta_{\rm SIF}$ due to larger secondary radii at the onset of CEE, allowing smaller values of $\beta$ to form a CBD. BH companions are the most massive, while giant progenitors of NSs are the least massive. Therefore, the BH companion does not need to spiral deep inside the NS progenitor to eject the envelope, and NS-BH binaries have the largest values of $\eta_{\rm SIF}$; these require larger values of $\beta$ to form CBDs.
4 Properties of post-common envelope circumbinary disk progenitors
@@
Here, we explore the pre and post-CEE characteristics of binary systems that result in DCOs and compare binaries that are likely to form CBDs after the termination of CEE with binaries that are not.
@@
In Figures 2-4, we present the properties of NS-NS for $\beta=0.304$, BH-BH for $\beta=0.197$, and NS-BH for $\beta=0.498$ binaries, respectively. These values of $\beta$ give that about half of the systems have a post-CEE CBD (Figure 1). Orange bins represent the distributions for cases where the leftover flattened post-CEE envelope has enough angular momentum to form a CBD according to the toy model presented in Section 3.1. Blue bins are for binary systems that form no CBDs. The upper panels of each figure show DCO binaries that will merge within Hubble time by emitting gravitational waves. In contrast, in the lower panels, we show binaries where the compact objects are too far apart to merge. The factor $f_{\rm CCSN}$ displayed in the left insets of Figures 2-4 is the percentage of each sub-class of systems from all core-collapse supernovae in our sample.
@@
Figure 2: Properties of NS-NS binary systems for $\beta=0.304$ (see equation 2) with (orange bins) and without (blue bins) post-CEE CBDs. The histograms show the distributions of the orbital semi-major axis $a_{\rm i}$ (left panels), and the mass ratio $q\equiv M_{\rm CO}/M_{\rm*,i}$ (middle panels) at the onset of CEE, and of the orbital separation after CEE divided by the radius of the giant at the onset of CEE, $\eta_{\rm SIF}\equiv a_{\rm f}/R_{\rm*,i}$. The vertical axis is the percentage of each column from the total number of systems in the plotted sub-population. The upper row displays the distributions of systems that will merge within Hubble time and the bottom row presents distributions for systems that are not expected to merge. The factor $f_{\rm CCSN}$ is the fraction of the systems relative to all core-collapse supernovae.
Figure 3: Similar to Figure 2 for BH-BH binary systems with $\beta$=0.197 .
Figure 4: Similar to Figure 2 for NS-BH binary systems with $\beta$=0.498 .
@@
The lower panels of Figures 2-4 show that NS-NS and BH-BH binaries that will not merge in Hubble time do not form post-CEE CBDs, while $\simeq 35\%$ of non-merging NS-BH binaries form CBDs after CEE of the BH and the NS progenitor; these are systems that satisfy $\eta_{\rm SIF}=a_{\rm f}/R_{\rm*,i}\lesssim 0.01$ (right panel of Figure 4). For higher values of $\beta$ we find non-merging NS-NS and BH-BH binaries in this regime as well. However, a substantial fraction of these non-merging binaries are expected to form CBDs only if the majority of DCO systems do so.
In the left panel of Figures 2-4 we show the semi-major axis at the onset of the CEE, $a_{\rm i}$. While for NS-NS binaries (Figure 2) and BH-BH binaries (Figure 3) $a_{\rm i}$ does not completely separate systems that form post-CEE CBDs from those who do not, Figure 4 shows that there is a clear distinction in $a_{\rm i}$ between two groups of NS-BH binaries; those that form CBDs have larger values of $a_{\rm i}$.
The progenitors of NS-NS and BH-BH binaries tend to have small initial mass ratios, in contrast to the case of NS-BH binaries. At the onset of the CEE, a relatively massive BH might enter the envelope of the NS’s progenitor, resulting in binary systems which are closer in mass. In the middle panel of Figures 2-4, we present the binary mass ratio at the beginning of the CEE stage. Comparing the three figures, we see indeed that the NS-BH progenitors have the largest ratio of the compact companion mass $M_{\rm CO}$ to the giant’s mass at the onset of the CEE, $M_{\ast,i}$, which can explain their different behavior. We also note that in all cases, DCO binaries that form CBDs have more extreme mass ratios (smaller values of $q\equiv M_{\rm CO}/M_{\ast,i}$) as expected from equation 4. A larger difference between the components of the binary is also coincident with stronger interaction between the surviving binary and the CBD, which is expected to shrink the binary in most cases (e.g., Valli et al. 2024).
The right panels of figures 2-4 show that the in-spiral factor, $\eta_{\rm SIF}\equiv a_{\rm f}/R_{\rm*,i}$, best distinguishes between systems that form or not CBDs in all DCO binaries. On average, the smaller this ratio is, the larger the ratio of the initial system’s specific angular momentum to the one required to form a CBD, hence favoring its formation. The values of $\eta_{\rm SIF}$ are notably the smallest for BH-BH binaries (right panel of Figure 3) and the largest for NS-BH binaries (right panel of Figure 4), which is coincidence with the difficulties to form CBDs according to the values of $\beta$ we found in Figure 1.
5 Implications for double compact object binaries and transients
@@
When an NS or a BH are close enough to a giant star, they can engage in a CEE phase where the compact object unbinds the envelope gas as it spirals in towards the center of the giant star. In cases where the entire envelope is ejected, the stripped core and the compact object remain in a contact binary, where the core could explode as a stripped-envelope supernova (e.g., Laplace et al. 2021; Vartanyan et al. 2021) forming another compact object. If the natal kick from the supernova is low enough, the binary system remains bound, resulting in a DCO binary. However, forming a post-CEE CBD encompassing the core and the compact object and its interaction with the binary system might affect the orbital evolution. If the binary system loses angular momentum to the CBD through viscous torques, the orbital separation will shrink, altering the formation and merger rates of DCO binaries in several ways.
@@
In cases where the compact object and the core are relatively close at the end of CEE, the interaction with the CBD could lead to a merger between the core and the compact object (e,g., Wei et al. 2024), giving rise to luminous transients such as fast blue optical transients and long gamma-ray bursts (e.g., Soker et al. 2019; Soker 2022; Metzger 2022  [^4] We note that the CBDs studied in Metzger (2022) to account for fast blue optical transients are much wider than the CBDs we explore here.; Grichener 2024), hence reducing the number of system that continue their evolution towards DCO binaries. Figure 5 shows the predicted periods of binary systems that form post-CEE CBDs and hence have the potential to trigger these transients as a function of the total binary mass at the onset of CEE.
@@
Figure 5: Orbital period vs total mass of the binary systems at the onset of CEE for progenitors of NS-NS binaries (blue dots), BH-BH binaries (green dots), and NS-BH binaries (red dots) that form a post-CEE CBD. Above and to the right: histograms showing the distributions of both quantities for the different DCO binaries subclasses. Each panel separately sums to $100\%$
@@
If the stripped core and the compact object begin further apart, however, then a compact object-core merger is prevented and the system would still evolve towards a DCO binary. The reduction in the orbital separation could lead to DCO merger in Hubble time and gravitational wave emission in systems that would have been too far to merge without the interaction with the CBD (e.g., Wei et al. 2024; Vetter et al. 2024). Such a scenario might also result in detectable long gamma-ray burst-like transient (Janiuk et al., 2013).
@@
In the left panel of Figure 6, we show the percentage of systems that result in NS-NS, BH-BH, and NS-BH binary systems in our stellar population, namely, all binary stars with primary initial mass in the range $5\leq M_{\rm 1,ZAMS}\leq 100$ The green columns represent DCO binaries that evolved through a CEE phase between the first compact object and the giant star. In contrast, the purple columns indicate systems formed via alternative evolutionary channels. We distinguish between systems in which the binary is expected to merge within Hubble time (middle panel) and in which the stars are not close enough to merge (right panel). While most NS-NS and NS-BH systems involve a giant-compact object CEE, this is not the case for BH-BH binaries, where most systems do not experience this evolutionary phase (as seen in the left panel). In all DCOs that merge, however, CEE is essential to decrease the orbital separation (middle panel; similarly to e.g., Mapelli et al. 2017; Vigna-Gómez et al. 2018; Giacobbo & Mapelli 2018; Chruslinska et al. 2018; Neijssel et al. 2019; Broekgaarden et al. 2021; see also Mandel & Broekgaarden 2022 and references therein).
@@
Figure 6: Percentage of binary systems that evolve to DCO binaries from the overall number systems in our binary stellar population, i.e., all primary stars in the range $5\leq M_{\rm 1,ZAMS}\leq 100$. The left panel shows all DCO binaries that went through a CEE phase between the compact objects and the giant star throughout their evolution (purple columns; See Figure 2 in Grichener 2023), or that evolved via other channels (green columns). The middle and right panels present the same information for DCOs that merge within Hubble time and for DCO binaries that are too wide to merge, respectively.
@@
Let us analyze each class of DCO binaries separately. The vast majority of NS-NS binary systems, whether they are expected to merge or not, are formed through an evolutionary pathway that includes CEE between an NS and a giant star (Figure 6). Since the Keplerian specific angular momentum around the post-CEE binary systems that would evolve to non-merging NS-NS binaries is high compared to the angular momentum of the flattened envelope, a CBD is not expected to form in these systems (lower panel of Figure 2), and therefore cannot aid them to merge. In NS-NS binaries that are expected to merge, though, a CBD could form relatively easily (Figure 1), potentially triggering a post-CEE merger of the NS with the exposed core of the giant star. Therefore, we conclude that in the case of NS-NS binaries, forming a CBD after CEE could only lower the event rates of NS-NS mergers.
@@
We can see in Figure 6 that a large fraction of BH-BH binaries, and in particular most binaries that are not expected to merge (right panel), do not involve a CEE phase in their evolution, implying their overall formation rates will not be significantly affected by a potential post-CEE CBD. The lack of CBD formation in non-merging BH-BH binaries (lower panel of Figure 3) strengthens this conclusion and indicates that BH-BH merger rates are unlikely to increase when considering the formation of CBDs. For BH-BH binaries that merge, however, CEE is still the dominant formation channel. Considering a large fraction of these binaries are likely to form CBDs (Figure 1) that might shrink the orbit of the binary, in some cases BH-core mergers occur earlier in the evolution, reducing the number of BH-BH mergers.
@@
Many NS-BH binaries go through a CEE of the BH and the giant (Figure 6); in all these, the BH is the descent of the primary star and the giant of the secondary star. Many of these end with CBD according to our criteria, implying that the CBD-induced early merger of the BH with the giant’s core can reduce the number of NS-BH binaries. Contrary to NS-NS binaries, some post-CEE non-merging NS-BH systems can form a CBD (lower panel of Figure 4). Losing angular momentum to the CBD will bring the stars in the surviving binary closer, resulting either in the merger of the core with the BH or forming a tighter NS-BH binary system that could merge in Hubble time. We note that there is a lower fraction of post-CEE non-merging NS-BH binaries compared to NS-BH binaries that are expected to merge. Therefore, it is more probable that the overall number of NS-BH mergers will decrease because of CBD-induced core-BH mergers. Given the difficulty of NS-BH binaries to form a CBD (Figure 1), the merger rates will likely not change much.
@@
Another potential implication of the change in the orbital separation due to the interaction of the binary with the CBD would be changes in the predicted delay-time distributions of DCO mergers. If a sufficient number of NS-NS binaries contract, for instance, the steeper delay-time distribution (prompter mergers) might be able to contribute early enough NS-NS mergers to explain some of the heavy elements formation in the early Universe, as required by recent studies (e.g., Maoz & Nakar 2024; Beniamini & Piran 2024).
6 Summary and discussion

### v2

@@ III.2 The role of the parameters $\beta$ and $\zeta$
To demonstrate the role of $\beta$ in determining the number of post-CEE systems that form a CBD, we re-analyze the population synthesis data from Grichener (2023) and examine three final populations that go through the CEE of a giant star with a compact object, and end as NS-NS, BH-BH, and NS-BH binaries.
In Figure 1, we present the percentage of NS-NS systems (blue dots), BH-BH systems (green dots), and NS-BH systems (red dots) that form a post-CEE CBD as a function of the parameter $\beta$. As expected, higher values of $\beta$, which represent systems that retain more of their initial angular momentum, correspond to a greater percentage of systems forming a post-CEE CBD. For BH-BH binaries and NS-BH binaries we note a sharp transition in $\beta$ between systems that are likely to form CBDs and those where such disks are not expected, while for NS-NS binaries the dependence is weaker. While recent papers (e.g., Izzard and Jermyn 2023) show a preference towards high values of $\beta$, due to the large uncertainties involved we show the results of a case where half of the systems will form CBDs. When varying the value of $\beta$, we find that qualitative trends in our results would change only if most systems formed CBDs. The graph shows that a CBD is formed in about half of the systems for $\beta=0.656$ in BH-BH binaries, $\beta=1.859$ in NS-NS binaries, and $\beta=0.713$ in NS-BH binaries.
We note from equations 4 and 5 that the parameter $\zeta$, which accounts for the mass distribution of the CBD, scales linearly with $\beta$. This implies that adopting a different mass distribution (i.e., a different $\zeta$) rescales the value of $\beta$ required to produce a fixed number of CBDs in each sub-population by the same multiplicative factor. Assuming a thin ring that viscously spreads to a CBD, corresponding to the absolute minimum mean specific angular momentum required for CBD formation, for instance, is represented by $\zeta=1$, and yields $\beta$ values that are a factor of four smaller than those quoted above for $\zeta=4$.
Figure 1: The percentage of NS-NS (blue dots), BH-BH (green dots), and NS-BH (red dots) systems that can form a CBD, from all systems that go through NS/BH-giant CEE phase and end as NS-NS systems, BH-BH systems, and NS-BH systems, respectively, taking a more restrictive condition for CBD formation ($\zeta=4$). Assuming the minimum specific angular momentum needed to form a thin ring that can subsequently viscously spread into a CBD (i.e., $\zeta=1$), the corresponding $\beta$-axis should be rescaled downwards by a factor of four.
IV Properties of post-common envelope circumbinary disk progenitors
@@
Here, we explore the pre and post mass transfer characteristics of binary systems that result in DCOs and compare binaries that are likely to form CBDs after the termination of CEE with binaries that are not.
@@
In Figures 2-4, we present the properties of NS-NS for $\beta=1.859$, BH-BH for $\beta=0.656$, and NS-BH for $\beta=0.713$ binaries, respectively. These values of $\beta$ indicate that about half of the systems have a post-CEE CBD (Figure 1). Orange bins represent the distributions for cases where the leftover flattened post-CEE envelope has enough angular momentum to form a CBD according to the toy model presented in Section III.1. Blue bins are for binary systems that form no CBDs. The upper panels of each figure show DCO binaries that will merge within Hubble time by emitting gravitational waves. In contrast, in the lower panels, we show binaries where the compact objects are too far apart to merge. The factor $f_{\rm CCSN}$ displayed in the left insets of Figures 2-4 is the percentage of each sub-class of systems from all core-collapse supernovae in our sample, which we present for observational reasons. The interaction of the CBD with the binary system might lead to a luminous transient triggered by the merger of the NS or the BH with the exposed core (see section V for a discussion). This transient event is termed common envelope jets supernova (CEJSNe; e.g., Soker and Gilkis 2018). At an earlier stage, when the NS/BH accretes mass in the envelope and launches jets, the bright transient is called a CEJSN impostor (e.g., Gilkis et al. 2019). Initially, a CEJSN or a CEJSN impostor might be misclassified as a CCSN. Later peculiar properties, e.g., no radioactive products (for CEJSN impostors), longer activity, and larger total energy than typical CCSNe, make the differentiation.
@@
Figure 2: Properties of NS-NS binary systems for $\beta=1.859$ (see equation 4) with (orange bins) and without (blue bins) post-CEE CBDs. The histograms show the distributions of the orbital semi-major axis $a_{\rm RL}$ (left panels), and the mass ratio $q\equiv M_{\mathrm{CO},RL}/M_{\rm*,RL}$ (middle panels) at the onset of RLOF, and of the orbital separation after CEE $a_{\rm f}$. The vertical axis represents the number of systems per bin, normalized such that the total sum over all bins is equal to one. The upper row displays the distributions of systems that will merge within Hubble time and the bottom row presents distributions for systems that are not expected to merge. The factor $f_{\rm CCSN}$ is the fraction of the systems relative to all core-collapse supernovae.
Figure 3: Similar to Figure 2 for BH-BH binary systems with $\beta$=0.656 .
Figure 4: Similar to Figure 2 for NS-BH binary systems with $\beta$=0.713 .
@@
The lower panels of Figures 2-4 show that most NS-NS and BH-BH binaries that will not merge in Hubble time do not form post-CEE CBDs, while $\simeq 25\%$ of non-merging NS-BH binaries form CBDs after CEE of the BH and the NS progenitor. By definition, for a high value of $\beta$ all double compact objects will form CBDs, including non-merging NS-NS and BH-BH binaries. However, we find that it is less likely for a substantial fraction of non-merging DCOs to form CBDs, since it would require high values of $\beta$.
In the left and middle panels of Figures 2-4 we show the semi-major axis and mass ratio of the binary systems at the onset of RLOF, respectively. The progenitors of NS-NS binaries tend to have smaller initial orbital separations and mass ratios, in contrast to the case of BH-BH binaries and NS-BH binaries, which might be a source of difference in the results. However, despite the fact that we would intuitively think that these two separations would play a strong role in determining the angular momentum reservoir available for CBD formation, they do not separate systems that form post-CEE CBDs from those who do not. This is due to the non-trivial dependence of the angular momentum on the different masses involved in the problem. The right panels of figures 2-4 show that the final orbital separation of the binary, that determines that boundaries of the disk in case it forms (Section III.1), is not a determining factor in CBD formation as well.
V Implications for double compact object binaries and transients
@@
When an NS or a BH is close enough to a giant star, they can engage in a CEE phase where the compact object unbinds the envelope gas as it spirals in towards the center of the giant star. In cases where the entire envelope is ejected, the stripped core and the compact object remain in a contact binary, where the core could explode as a stripped-envelope supernova (e.g., Podsiadlowski et al. 1992; Laplace et al. 2021; Vartanyan et al. 2021; Lu et al. 2025), forming another compact object. If the natal kick from the supernova is low enough, the binary system remains bound, resulting in a DCO binary. However, forming a post-CEE CBD encompassing the core and the compact object and its interaction with the binary system might affect the orbital evolution. If the binary system loses angular momentum to the CBD through gravitational torques, the orbital separation will shrink, altering the formation and merger rates of DCO binaries in several ways.
@@
In cases where the compact object and the core are relatively close at the end of CEE, the interaction with the CBD could lead to a merger between the core and the compact object (e,g., Wei et al. 2024), giving rise to luminous transients, CEJSNe, appearing as, e.g., fast blue optical transients and long gamma-ray bursts (e.g., Soker et al. 2019; Soker 2022; Metzger 2022  [^5] We note that the CBDs studied in Metzger (2022) to account for fast blue optical transients are much wider than the CBDs we explore here.; Grichener 2025), reducing the number of systems that continue evolving towards DCO binaries. Figure 5 shows the predicted periods of binary systems that form post-CEE CBDs and hence have the potential to trigger these transients as a function of the total binary mass at the onset of CEE. Since CEJSNe might be classified as peculiar CCSNe, at least at early times of the event, it is important to present the properties of their progenitors for future identifications in observations. Figure 5 further emphasizes the properties of the systems we study, which can be used to re-analyze peculiar CCSNe and reclassify them as CEJSN events.
@@
Figure 5: Orbital period vs total mass of the binary systems at the onset of CEE for progenitors of NS-NS binaries (blue dots), BH-BH binaries (green dots), and NS-BH binaries (red dots) that form a post-CEE CBD. Above and to the right: histograms showing the distributions of both quantities for the different DCO binaries subclasses. We normalized the histograms such that in each sub-panel the sum over the bins is equal to one.
@@
If the stripped core and the compact object begin further apart, however, then a compact object-core merger is prevented and the system would still evolve towards a DCO binary. The reduction in the orbital separation could lead to DCO merger in Hubble time and gravitational wave emission in systems that would have been too far to merge without the interaction with the CBD (e.g., Wei et al. 2024; Siwek et al. 2023 Vetter et al. 2024). In the case of a BH-BH binary system, the merger might result in detectable long gamma-ray burst (Janiuk et al., 2013).
@@
In the left panel of Figure 6, we show the percentage of systems that result in NS-NS, BH-BH, and NS-BH binary systems in our stellar population, namely, all binary stars with primary initial mass in the range $5\leq M_{\rm 1,ZAMS}\leq 100$ . The green columns represent DCO binaries that evolved through a CEE phase between the first compact object and the giant star. In contrast, the purple columns indicate systems formed via alternative evolutionary channels. We distinguish between systems in which the binary is expected to merge within Hubble time (middle panel) and in which the stars are not close enough to merge (right panel). While most NS-NS and NS-BH systems involve a giant-compact object CEE, this is not the case for BH-BH binaries, where most systems do not experience this evolutionary phase (as seen in the left panel). In the majority of DCOs that merge, however, CEE is essential to decrease the orbital separation (middle panel; similarly to other population synthesis studies like e.g., Mapelli et al. 2017; Vigna-Gómez et al. 2018; Giacobbo and Mapelli 2018; Chruslinska et al. 2018; Neijssel et al. 2019; Broekgaarden et al. 2021; Mandel and Broekgaarden 2022 and references therein; see Gallegos-Garcia et al. 2021 for a different conclusion drawn based on detailed stellar evolution simulations).
@@
Figure 6: Percentage of binary systems that evolve to DCO binaries from the overall number systems in our binary stellar population, i.e., all primary stars in the range $5\leq M_{\rm 1,ZAMS}\leq 100$. The left panel shows all DCO binaries that went through a CEE phase between the compact objects and the giant star throughout their evolution (green columns; See Figure 2 in Grichener 2023), or that evolved via other channels (purple columns). The middle and right panels present the same information for DCOs that merge within Hubble time and for DCO binaries that are too wide to merge, respectively.
@@
Let us analyze each class of DCO binaries separately. The vast majority of NS-NS binary systems, whether they are expected to merge or not, are formed through an evolutionary pathway that includes CEE between an NS and a giant star (Figure 6). Since the Keplerian specific angular momentum at the expected inner radius of the disk in post-CEE binary systems that would evolve to non-merging NS-NS binaries is high compared to the angular momentum of the flattened envelope, a CBD is not expected to form in these systems (lower panel of Figure 2), and therefore cannot aid them to merge. In NS-NS binaries that are expected to merge, though, a CBD could form relatively easily (Figure 1), potentially triggering a post-CEE merger of the NS with the exposed core of the giant star. Therefore, we conclude that in the case of NS-NS binaries, forming a CBD after CEE could only lower the event rates of NS-NS mergers.
@@
We can see in Figure 6 that a large fraction of BH-BH binaries, and in particular most binaries that are not expected to merge (right panel), do not involve a CEE phase in their evolution, implying their overall formation rates will not be significantly affected by a potential post-CEE CBD. The lack of CBD formation in non-merging BH-BH binaries (lower panel of Figure 3) strengthens this conclusion and indicates that BH-BH merger rates are unlikely to increase when considering the formation of CBDs. For BH-BH binaries that merge, however, CEE is still the dominant formation channel according to our population synthesis study (see Neijssel et al. 2019 and Gallegos-Garcia et al. 2021 for another view). Considering a large fraction of these binaries are likely to form CBDs (Figure 1) that might shrink the orbit of the binary, in some cases BH-core mergers occur earlier in the evolution, reducing the number of BH-BH mergers. Figure 7 shows that overall the same conclusion holds for a metallicity of $Z=0.01Z_{\rm\odot}$ which is more typical for merging binary black holes, with the CEE being less prominent but with an higher overall number of binary black holes that evolve through CEE and result in DCO mergers (see section VI for a discussion of the effects of metalicity on our results).
@@
Many NS-BH binaries go through a CEE of the BH and the giant (Figure 6); in all these, the BH is the descendant of the primary star and the giant of the secondary star. Many of these end with CBD according to our criteria, implying that the CBD-induced early merger of the BH with the giant’s core can reduce the number of NS-BH binaries. Contrary to NS-NS binaries, some post-CEE non-merging NS-BH systems can form a CBD (lower panel of Figure 4). Losing angular momentum to the CBD will bring the stars in the surviving binary closer, resulting either in the merger of the core with the BH or forming a tighter NS-BH binary system that could merge in Hubble time. We note that there is a lower fraction of post-CEE non-merging NS-BH binaries compared to NS-BH binaries that are expected to merge. Therefore, it is more probable that the overall number of NS-BH mergers will decrease because of CBD-induced core-BH mergers. However due to contradictory effects, the merger rates will likely not change much.
@@
Another potential implication of the change in the orbital separation due to the interaction of the binary with the CBD would be changes in the predicted delay-time distributions of DCO mergers. If a sufficient number of NS-NS binaries contract, for instance, the steeper delay-time distribution (prompter mergers) might be able to contribute early enough NS-NS mergers to explain some of the heavy elements formation in the early Universe, as required by recent studies (e.g., Maoz and Nakar 2024; Beniamini and Piran 2024). This is the subject of a future study
Figure 7: Percentage of binary systems that evolve to BH-BH binaries from the overall number systems in our binary stellar population for a metallicity of $Z=0.01Z_{\rm\odot}$. Left panel: all BH-BH binaries that evolved through a CEE phase between the BH and the giant star (green columns), or that evolved via other channels (purple columns). Half of the BH-BH binaries form a post-CEE CBD (orange columns) for $\beta=0.709$. Middle panel: same as left panel for BH-BH binaries that merge within Hubble time. Right panel: same as left panel for systems that are not expected to merge.
VI Summary and discussion